# Perfect codes over non-prime power alphabets: an approach based on Diophantine equations

### v1

@@ 5. Concluding remarks
which has conductor $N=16328471028588=2^{2}\cdot 3^{3}\cdot 47^{2}\cdot 8273^{2}$. We are able to show that $E_{2\cdot 47^{2}}$ has rank $1$, but the conductor is too large to apply the Heegner point methodology and, consequently, we are unable to find a generator of $E_{2\cdot 47^{2}}(\mathbb{Q})$.
More generally, our techniques will not be able to resolve the problem if $q$ has large prime divisors. For this reason, we have been able to solve many cases of $q>200$ by assuming that all prime divisors of $q$ lied in the set $\{2,3,5,7,11\}$. Even on this case, the necessary computational effort is significant, and so it is unlikely that this could be generalised to larger values of $q$.
Apart from the computational challenges, we remark that this methodology is not able to deal with families of $q$, but only with individual values. In order to prove Conjecture 1, even just for $e=2$, it seems likely that significantly new ideas will need to be introduced.
@@
However, the methods that we present here could be used to approach perfect quantum codes, by mimicking this approach and changing the perfect code condition (1) by the perfect quantum code condition:
$M={q^{n}}\left/\sum_{j=0}^{e}{\binom{n}{j}}(q^{2}-1)^{j}\right..$
@@
Li and Xing [16] classified perfect quantum codes over $q-$dimensional qudits, where $q$ is a prime power. In addition, they proved an analogue of Lloyd’s theorem for quantum codes. This makes the situation completely analogous to the classical case and, for this reason, it seems plausible that the approach that we presented here would directly generalise to quantum codes.

### v2

@@ 5.1. Limitations of our methodology
which has conductor $N=16328471028588=2^{2}\cdot 3^{3}\cdot 47^{2}\cdot 8273^{2}$. While we are able to show that $E_{2\cdot 47^{2}}$ has rank $1$, $N$ is too large to apply the Heegner point methodology and, consequently, we are unable to find a generator of $E_{2\cdot 47^{2}}(\mathbb{Q})$.
If $q$ has a large prime divisor $p$, there is a slightly different problem. In this instance, and for $q$ large, we need to use the Magma IntegralPoints subroutine to find integral points on the Mordell curve (21). However, this relies on the use of linear forms in $p-$adic elliptic logarithms (see [28]) and the bounds obtained by applying this method are astronomical as $p$ grows, which makes it impossible to find all solutions to (13). In fact, this is the main reason why we needed to exclude prime divisors larger than $13$ if $q\leq 600$.
Due to the high computational burden, we believe that the results that we present here cannot be significantly extended without the introduction of new algorithms to resolve the Ramanujan–Nagell equation (13). Promising results in this direction are the newly-improved estimates for linear forms in three logarithms by Mignotte and Voutier [22] and some work on the generalised Ramanujan–Nagell equation by Mutlu, Le and Soydan [24], which, if extended, could allow for a more efficient resolution of (13).
5.2. Future lines of work
In this paper, we used extensively two features of perfect $2-$error correcting codes: the perfect code condition (10) and the fact that Lloyd’s Theorem (Theorem 4) applies. As long as these two conditions are satisfied, it seems plausible that our approach could generalise.
@@ 5.2. Future lines of work
For instance, the methods that we present here could be used to approach perfect $2-$error correcting quantum codes, by mimicking this approach and changing the perfect code condition (5) by the perfect quantum code condition:
(31) $M={q^{n}}\left/\sum_{j=0}^{e}{\binom{n}{j}}(q^{2}-1)^{j}\right..$
@@
Li and Xing [19] classified perfect quantum codes over $q-$dimensional qudits, where $q$ is a prime power. In addition, they proved an analogue of Lloyd’s theorem for quantum codes. This makes the situation completely analogous to the classical case and, for this reason, it seems reasonable that the approach that we presented here would directly generalise to quantum codes.
6. Conclusions
The main findings of this paper are Theorems 2 and 3, where we show that perfect $2$-error correcting codes over $q-$ary alphabets do not exist for more than $170$ new values of $q$, and show that, for a fixed value of $q$, there can only be finitely many $2-$error correcting perfect codes.
In previously existing literature, no values of $q$ with more than three prime factors were considered and, in addition, the methods used were specific to the cases at hand and could not be generalised. In addition, there had been no new results in over $50$ years, showing a clear gap in the literature.
In this paper, we successfully address this gap by showing non-existence of perfect codes for some values of $q$ with four prime factors ($q=210,330,462$) and provide a method which would theoretically apply to any fixed value of $q$, with the only existing problems arising from current computational technology. Our results improve the number of $q$ for which $2-$error correcting codes have been classify by more than 2,000%.
We remark that these results are also relevant from a practical point of view, since they show that, for many values of $q$, the optimal information transmission scheme while attempting to correct $2$ errors cannot be that of perfect codes and, consequently, alternative schemes need to be considered in order to achieve the best possible codes.
In order to verify all computations carried out in this paper, we provide the reader with the Magma code available in the author’s GitHub repository (accessible at https://github.com/PJCazorla/perfect-q-ary-codes), which can be used to check the correctness of Theorem 2.
Finally, we remark that our results are consistent with the existing literature in the cases where they overlap and, in addition, they support Conjecture 1, which is widely believed to be true.